OpenStudy (anonymous):

The odds against Lounging Larry winning the horse race are 5:12. Determine the probability that Lounging Larry wins the horse race.

OpenStudy (anonymous):

7:12

OpenStudy (anonymous):

unmath thats wrong

OpenStudy (anonymous):

its 5:17

OpenStudy (anonymous):

@unmathguy its cuz you are going from a odds to a probability

OpenStudy (anonymous):

wait it'll be 7:17

OpenStudy (anonymous):

aight! i haven't been in a stat class since JH which was like 5 yrs ago! srry!

Directrix (directrix):

Odds against are 5:12. Think of as 5 against and 12 for winning Probability = number of favorable outcomes divided by total number of outcomes. P(Larry wins) = 12 / (5 + 12) = 12/17